40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14645     Entry time: Fri May 31 15:55:16 2019     In reply to: 14643
Author: gautam 
Type: Update 
Category: ALS 
Subject: PSL + X beat restored 

Coupling into the fast axis of the fiber:

The PM couplers I bought require that the light is coupled to the fast axis. The Thorlabs part that Andrew ordered, and which Anjali was using for the MZ experiment, was the opposite configuration, and so the input coupler K6XS mount was rotated to accommodate this polarization. The HWP was also rotated to cut the power into the fiber. I undid these changes. Mode-matching is ~65% (2.42mW/3.70mW) which isn't stellar, but good enough. The PER is ~15dB (ratio of power in fast axis to slow axis is ~40), which I verified using another collimator at the output, and a PBS + two photodiodes. Again isn't stellar but good enough.

EX laser temperature adjustment:

Rana adjusted the temperature of the main laser to 30.61 C. According to the calibration, the EX laser temperature needed to be ~32.8 C. It was ~31.2 C. I made the change by rotating the dial on the front panel of the EX laser controller. Fine adjustment was done using the temperature slider on the ALS screen. With an offset of ~+610 counts, I found a beat at ~80 MHz.

First look at PM beamsplitters:

From my initial test, the beat amplitude was stable to my moving of the fibers yes. The NF1611 DC monitor reports 2.6 V DC with only the EX light, and 3.15 V DC with only the PSL light. So I should probably cut the PSL power a little to improve the contrast. Assuming the 10 kohm DC transimpedance spec can be believed, this means the expected signal level is 4*sqrt(260uA * 315uA)*700V/A ~0.8 Vpp, and I see ~0.9 Vpp, so roughly things add up (this is actually more consistent with an RF transimpedance of 800V/A, which is maybe not unreasonable). The RF amps for routing this signal to the delay line has been borrowed for the 2um frequency noise experiemnt - I will reacquire it today and check the ALS noise performance.

So overall, I am happy with the performance of the current iteration of the BeatMouth.

ELOG V3.1.3-